OpenStudy (anonymous):

Find the derivative of the function. y =2^(5-x^2)

OpenStudy (anonymous):

the derivative of \(b^x\) is \(b^x\ln(b)\) so by the chain rule the derivative of \(b^{f(x)}\) is \(\ln(b)b^{f(x)}f'(x)\)

OpenStudy (anonymous):

That is actually a really good explanation @satellite73!

OpenStudy (anonymous):

so ln(2)2^(5-x^2)-2x?

OpenStudy (anonymous):

@KeithAfasCalcLover thanks @amadosw it is times not minus

OpenStudy (anonymous):

(-2x)

OpenStudy (anonymous):

yeah, like that you might want to clean it up a bit and write \[-2\ln(2)x2^{5-x^2}\]